OpenStudy (knudsona):

I really need some help with a couple questions. I have studied this lesson multiple times, but still struggle to understand it. It has to do with linear inequalities. Medals will be rewarded!!

OpenStudy (knudsona):

Here's the first question in having trouble with.. Which of the following points lie in the solution set to the following system of inequalities? y (Less than or equal) x - 5 y (Greater than or equal) -x - 4

hartnn (hartnn):

I am sure you are given some points in choices/options. lets say you have a point 1,2 how to determine whether it "lies in the solution set " ? simple, just plug in x=1, y =2 in both of your inequalities. check whether both comes out to be true or not . \(y \le x-5 \\ 2 \le 1-5 \\ 2\le -2 \\True!\) \(y\ge -x-4 \\ 2\ge -1-4 \\ 2\ge -5 \\True!\) Since , BOTH are true, the point 1,2 will lie in the solution set makes sense ?

OpenStudy (knudsona):

Im starting to get it, so with the options they give me, I would fit them into the problems and see if they work?

hartnn (hartnn):

thats correct :)

OpenStudy (knudsona):

Okay, thankyou. Its just extremely hard for me to learn on virtual school. Haha, you worded it a little better than they did.

OpenStudy (knudsona):

WAIT. Another little question.

hartnn (hartnn):

sure

OpenStudy (knudsona):

The options it gives me are like (example) (5,2) Would i plug the 5 into the X of one of the problems and them plug the 2 into the Y of the other?

hartnn (hartnn):

so, x= 5, y =2 plug these in both of your inequalities

OpenStudy (knudsona):

So i would plug them into BOTH, and if they turn out to be correct, that's the answer?

hartnn (hartnn):

yes, if BOTH turn out to be true, than you got your point :)

hartnn (hartnn):

*then

OpenStudy (knudsona):

THANKYOU THANKYOU THANKYOUUUUU. You have helped me save my math grade!!!

hartnn (hartnn):

I am glad to hear that :) and you're most welcome ^_^